OpenStudy (anonymous):

Use the table (in comments) to complete each part. A:Copy the table. Extend the table by writing expressions for y when x =5,6 and 7. B:Simplify each expression for y in the table from part (A). What pattern do you notice when you simplify the expressions? C: Write an equation that relaxes x and y. Use your equation to find the value pf y when x=15.

OpenStudy (anonymous):



OpenStudy (azureilai):

So the generic trend is to add the next \[2^{n}\] number in the y. for example, 2^2 =4 2^3=8 2^4=? 2^5=? 2^6=? You will be adding the last three numbers I just gave you to solve one by one onto the table. When you are done, you can add like terms on all rows and start working on section B.

OpenStudy (anonymous):

ill try. Thank you :)

OpenStudy (azureilai):

no problem

OpenStudy (anonymous):

so are you multiplying 2 and x?

OpenStudy (anonymous):

the given numbers in x shall I say

OpenStudy (anonymous):

I'm really really bad at math.. hence I'm on here lol

OpenStudy (azureilai):

For B, I think for every x difference, the y has a difference of 2^n. I usually use a graphing calculator to find the lines and equations, so I am not that good at algebraically solving these.

OpenStudy (anonymous):

oh ok so should I bump the question?

OpenStudy (azureilai):

yeah go ahead

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay, Thank you so much for trying to help though :)

OpenStudy (azureilai):

no problem, I hope you eventually get it figured out

OpenStudy (anonymous):

thanks,ya me too!

OpenStudy (anonymous):

@Luigi0210 you think you could help?